OpenStudy (anonymous):

Need this explained please! Find the first six terms of the sequence. a1 = 6, an = 2 • an-1

OpenStudy (crashonce):

just substitute 'n' with your number

OpenStudy (anonymous):

so: an=2*6-1? @CrashOnce

OpenStudy (crashonce):

ya so a2=2x6 -1 therefore a2 = 11

OpenStudy (crashonce):

sorry thats a6

OpenStudy (anonymous):

can you just rewrite it for me please? This is super confusing online!

OpenStudy (anonymous):

@CrashOnce

OpenStudy (crashonce):

a2=2*2-1=3 a3=2*3-1=5 a4=2*4-1=7 i think its like this may be wrong

OpenStudy (anonymous):

@luigi0210

OpenStudy (anonymous):

I'll try but what would I do with the 6? @CrashOnce

OpenStudy (anonymous):

just want to clarify, is it \[a_n=2\times a_{n-1}?\]

OpenStudy (anonymous):

yes! I didn't know how to write it like that.

OpenStudy (anonymous):

it seems geometric sequence to me... where in the we have a common ratio between terms... and 2 is the common ratio between terms.... and the first term \(a_1=6\)

OpenStudy (anonymous):

so \(a_2=2 \times 6 = 12\)

OpenStudy (anonymous):

and \(a_3=2 \times 12=24\) and so on up to the 6th term...

OpenStudy (anonymous):

so a4 would be 2*24=48?

OpenStudy (anonymous):

yup...

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

no problem...